OpenStudy (anonymous):

If 4 people can paint 6 walls in 30 minutes. How many minutes will it take for 7 people to paint 9 walls??

OpenStudy (anonymous):

4 eople --- 6 walls --- 30 mins. 1 person - 6/4th of a wall----- 30mins. 1 person ---1 wall-----30*4/6 mins= 20 mins Terefre If 7 people paint 9 walls means each person paints 9/7 th of a wall. Time taken for that = 20 * 9 / 7 Minutes. = 25.7 mins

OpenStudy (anonymous):

why do you multiply 30*4/6???

OpenStudy (anonymous):

did you use direct variation???

OpenStudy (anonymous):

Well. Its unitary method. U can think of ilik that. But its basiclly - A guy does 1/2 the work in 5mins. So he'll do the entire thing in 5 * 2 = 10 mins. And Yes, You are correct. Its direct variation.